OpenStudy (anonymous):

A car costs $25,000 and depreciates in value by 9% per year. How much will the car be worth after 4 years? Round to hundredths.

OpenStudy (johnweldon1993):

So you have a simple interest problem here... Amount = P( 1 - r)^t t = time (in years) r = rate (percent in decimal) P = principal/original amount So Amount = 25000(1 - .09)^4 This comes out to what?